OpenStudy (anonymous):

solve x^4 - 34x^2 = -225

OpenStudy (anonymous):

x^4 - 34x^2 + 225 = 0

OpenStudy (anonymous):

Let x^2 be a a^2 - 34a + 225 = 0

OpenStudy (anonymous):

so it is no solution?

OpenStudy (anonymous):

no you use quadratic or factoring to solve a^2 - 34a + 225 = 0 first

OpenStudy (anonymous):

a = 9 a = 25 x^2 = 9 x = plus/ minus 3 x^2 = 25 x = plus/minus 5